The sales representative informs you that there are 56 houses for sale with two floor plans still available. Use x to represent floor plan one and y to represent floor plan two. Write an equation that illustrates the situation.
x + y = 56
b. The sales representative later indicates that there are three times as many homes available with the second floor plan than the first. Write an equation that illustrates this situation. Use the same variables you used in Part a.

x + 3y = 56

c. Use the equations from Parts a. and b. of this exercise as a system of equations. Use substitution to determine how many of each type of floor plan is available. Describe the steps you used to solve the problem.
d. What are the intercepts of the equation from Part a. of this problem? What are the intercepts from Part b. of this problem? Where would the lines intersect if you solved the system by graphing?

Answer by stanbon(75887)

The sales representative informs you that there are 56 houses for sale with two floor plans still available. Use x to represent floor plan one and y to represent floor plan two. Write an equation that illustrates the situation.
x + y = 56
b. The sales representative later indicates that there are three times as many homes available with the second floor plan than the first. Write an equation that illustrates this situation. Use the same variables you used in Part a.
---
y = 3x
c. Use the equations from Parts a. and b. of this exercise as a system of equations. Use substitution to determine how many of each type of floor plan is available. Describe the steps you used to solve the problem
---
Substitute for "y" and solve for "x":
x + 3x = 56
4x = 56
x = 14 (smaller number of floor plans)
---
Since y = 3x, y = 3*14 = 42 (larger number of floor plans)
--------------------------------------------------------------
d. What are the intercepts of the equation from Part a. of this problem?
x+y = 56
If x = 0, y = 56
If y = 0, x = 56
----------------------------
What are the intercepts from Part b. of this problem?
y = 3x
If x = 0, y = 0
If y = 0, x = 0
-----------------------------
Where would the lines intersect if you solved the system by graphing?
(14,42)
==============
Cheers,
Stan H.